# Can LLM-based Financial Investing Strategies Outperform the Market in Long Run?

### v3

@@ 1. Introduction
Large language models (LLMs) are increasingly used in financial decision-making, especially for generating investment actions such as Buy, Hold, or Sell (Ding et al., 2024a; Fatouros et al., 2025). These so-called LLM timing-based investing strategies leverage LLMs’ ability to interpret historical and real-time data to autonomously trade. From sentiment-driven trading (Zhang et al., 2024a) to sophisticated multi-agent systems (Yu et al., 2023; Zhang et al., 2024b), a growing body of work has explored the potential of LLMs as autonomous financial agents.
Backtesting is the standard method for assessing investment strategies, simulating them on historical data to evaluate profitability and robustness (Chan, 2021). However, current LLM investing research suffers from fragmented, underdeveloped evaluation practices. Most studies assess performance over short periods, on few stock symbols, and often omit code release, limiting reproducibility. As summarised in Table 1, several recent methods evaluate over under a year, with fewer than ten stocks, and benchmark only against naïve baselines like Buy-and-Hold. Such short horizons and narrow stock universes lead to three well-documented sources of bias: survivorship bias (Garcia and Gould, 1993), where delisted or failed stocks are omitted; look-ahead bias (Chan, 2021), where future information inadvertently influences past decisions; and data-snooping bias (Bailey et al., 2015), where strategy performance is inflated through repeated testing on the same data. These biases can result in misleading performance assessments and undermine the validity of claimed improvements over traditional methods. This raises a central question: Can LLM-based investing strategies survive longer and broader robustness evaluations?
@@
While recent efforts such as Wang et al. (2025) and Hu et al. (2025) have addressed benchmarking for deep learning (DL)-based trading and LLM-based time-series forecasting, comprehensive evaluation of LLM-based investing strategies remains unaddressed. Separately, FinBen (Xie et al., 2024) provides a thorough FinLLM benchmark covering multiple tasks, including decision-making. However, as a broad FinLLM benchmark, FinBen’s backtesting still relies on a limited, hand-picked symbol set, which contains the aforementioned biases and lacks a professional backtesting pipeline or systematic comparison with traditional strategies. To fill this gap, we introduce FINSABER, a comprehensive framework for benchmarking LLM timing-based investing strategies that supports longer backtesting periods, a broader and more diverse symbol universe, and explicit bias mitigation. Specifically, our main contributions are:
@@ 2. Related Works
Recent work using LLMs as investors directly employ LLMs to make investing decisions (Ding et al., 2024a). The most common approach leverages LLMs’ sentiment analysis capabilities, using either general-purpose LLMs (e.g., GPT, LLaMA (Touvron et al., 2023), Qwen (Bai et al., 2023)) or fine-tuned financial variants like FinGPT (Yang et al., 2023) to generate sentiment scores for trading decisions (Lopez-Lira and Tang, 2023; Wu, 2024; Kirtac and Germano, 2024; Zhang et al., 2024a). However, these approaches stop short of forming complete trading strategies, which require not only directional forecasts, but also realistic liquidity sizing for mitigating impact, development of execution rules for trade timing and risk management, and incorporation of trading costs.
More advanced approaches move beyond sentiment scores by summarising and reasoning over multi-source financial text. For example, Fatouros et al. (2024) introduce a memory module that stores summarised financial data, retrieved during trading to guide decisions. Similarly, LLMFactor (Wang et al., 2024) learns to extract profitable factors from historical news aligned with price movements and applies them to future market forecasts.
A growing body of work incorporates LLM-based agents (Guo et al., 2024), where either one specialised agent or multiple collaborative agents are employed to perform financial analysis or predictions. Notable examples include FinMem (Yu et al., 2023), FinAgent (Zhang et al., 2024b), FinRobot (Yang et al., 2024), TradExpert (Ding et al., 2024b), FinCon (Yu et al., 2024), TradingAgents (Xiao et al., 2024) and MarketSenseAI 2.0 (Fatouros et al., 2025). Some models also incorporate reinforcement learning (RL) for iterative self-improvement (Ding et al., 2023; Koa et al., 2024).
@@ 4. Biases and Robustness Challenges in Backtesting LLM Investors
Robust evaluation of financial strategies demands carefully designed backtests. Unlike typical machine learning tasks with large, clean datasets, financial data is noisy, nonstationary, and limited in scope. As a result, backtests are especially prone to three major sources of bias: survivorship bias, look-ahead bias, and data-snooping bias, each of which can inflate perceived performance and lead to misleading conclusions (Chan, 2021).
@@
This occurs when backtests include only currently active stocks while ignoring delisted or bankrupt assets. Such omissions systematically overstate returns and understate risk (Joubert et al., 2024). A common cause is using today’s S&P 500 constituents as the historical investment universe. This practice introduces what Garcia and Gould (1993) call “preinclusion bias”, also a form of look-ahead bias where future index membership influences past decisions. The impact is well-documented: Grinblatt and Titman (1989) and Elton et al. (1996) estimate annual return distortions between 0.1% and 0.9%, and Brown et al. (1992) show that even small distortions can misrepresent performance persistence.
@@
Look-ahead bias arises when a strategy uses information that would not have been known at the time of decision-making (Chan, 2021). This includes selecting features, parameters, or symbols based on full-period outcomes, thereby introducing future knowledge into the backtest.
@@
Also known as multiple testing bias, this occurs when repeated experimentation on the same dataset leads to overfitting. In finance, where sample sizes are small and the signal-to-noise ratio is very low, this bias is particularly problematic. Bailey et al. (2015) showed that evaluating strategies on overlapping data inflates false positive rates, and that standard hold-out validation techniques often fail to guard against this issue.
@@
Addressing these biases requires evaluating strategies across longer periods and broader asset universes. For daily trading, at least three years of data is generally recommended, while weekly and monthly strategies benefit from 10 to 20 years or more (Bailey et al., 2015). Gatev et al. (2006) tested pairs trading on 40 years of daily data, but Do and Faff (2010) extended this to 48 years and found profitability declined, highlighting the need for long-term evaluation. Likewise, recent deep learning models in finance rely on multi-year datasets to ensure robustness (Feng et al., 2019; Wang et al., 2022; Xia et al., 2024).
@@ 5. FINSABER
We incorporate a comprehensive collection of strategies across multiple paradigms to ensure robust benchmarking. The timing-based strategies include open-source LLM investors (FinMem (Yu et al., 2023), FinAgent (Zhang et al., 2024b)), traditional rule-based approaches (Buy and Hold, Moving Average Crossover, Bollinger Bands (Bollinger, 2002), Trend Following (Wilcox et al., 2005)), ML/DL forecaster-based methods (ARIMA, XGBoost), and RL-based strategies (A2C, PPO, TD3, SAC implemented via FinRL (Liu et al., 2022) framework). Selection-based strategies encompass random K selection, Momentum Factor Selection (based on past returns), Volatility Effect Selection (selecting low-volatility stocks), and the stocks selection agent from the FinCon (Yu et al., 2024) framework. This diverse strategy base enables comprehensive performance comparison across different methodological approaches while maintaining extensibility for custom implementations. More technical details of the strategies are available in Appendix B.
@@
High returns alone do not imply strategy quality. Risk-adjusted metrics such as SPR and STR are more informative, especially in finance where capital efficiency and downside risk are critical (Chan, 2021). These metrics are standard in the literature (Cont, 2001; DeMiguel et al., 2007) and widely used in recent LLM-based investing benchmarks (Zhang et al., 2024b; Yu et al., 2024). Formal definitions and formulas are provided in Appendix C.
@@ 6.2. Fair Comparisons with the Composite Approach
To overcome the aforementioned biases, we introduce the Composite evaluation setup within FINSABER. This setup integrates systematic selection-based strategies to expand and diversify the stock universe, explicitly addressing survivorship and data-snooping biases. Specifically, we use four unbiased stock selection approaches from the strategies base (details in Appendix B): Random Five, Momentum Factor (Muller and and, 2010), Volatility Effect (Blitz and Vliet, 2007), and the FinCon Selection Agent in the FinCon (Yu et al., 2024) framework.
@@ 6.3. Statistical Validation and Behavioural Diagnostics of LLM Agents
To validate our findings from the composite backtests and diagnose the underlying drivers of LLM agent performance, we conduct a unified statistical and behavioural analysis. First, we conduct paired t-tests comparing Buy and Hold, FinMem, and FinAgent across both Selected 4 (Table 3) and Composite (Table 4) setups. Second, we dissect the agents’ behavioural characteristics by examining their drawdown profiles, alpha ($\alpha$) and beta ($\beta$) decomposition, and trading turnover across the different selection environments. These metrics are obtained by regressing the strategy’s excess returns against the market’s excess returns based on the Capital Asset Pricing Model (CAPM) (Sharpe, 1964). The model is defined as: $R_{s}-R_{f}=\alpha+\beta(R_{m}-R_{f})+\epsilon$, where $R_{s}$ is the return of the strategy, $R_{m}$ is the market return, $R_{f}$ is the risk-free rate, and $\epsilon$ is the idiosyncratic residual. In this model, $\beta$ measures the strategy’s systematic risk or volatility relative to the market, while $\alpha$ represents the portion of the return not explained by market exposure, often considered a measure of strategy-specific skill.
@@
Table 5 reports t-tests and p-values for the previous results, testing the null hypothesis of equal performance distributions. Under the selective period, statistical significance is inconsistent and limited mostly to individual stocks. However, after mitigating biases through the composite setup, the p-values drop substantially, indicating the market baseline (Buy and Hold) significantly outperforms both LLM strategies across all robust setups. Notably, while FinAgent tends to outperform FinMem when biases are controlled, both still significantly underperform simple market baselines. Furthermore, the behavioural analysis in Table 6 reveals that this underperformance is rooted in a lack of genuine skill; neither LLM agent generates statistically significant alpha, with all measured p-values exceeding 0.34. This finding robustly supports our main thesis that the claimed superiority of these models does not hold under rigorous evaluation, aligning with the Efficient Market Hypothesis (Lo, 2004).
@@ 7. Market Regime Analysis
Another key question in evaluating LLM-based investing strategies is whether they adapt appropriately across varying market conditions. Financial markets exhibit time-varying predictability and uncertainty across different economic, financial, and political regimes (Kim et al., 2011). Some strategies may exploit these variations, while others may struggle to adapt. Distinct market environments—bull, bear, and sideways—present unique challenges and opportunities: bull markets reward aggressive positioning and high exposure, bear markets require effective risk management, and sideways markets test a strategy’s ability to navigate uncertainty in the absence of clear trends. By decomposing performance across these regimes, it is possible to determine whether strategies are overly conservative and miss opportunities during bullish periods, or excessively aggressive and incur significant losses during downturns. Understanding these regime-specific behaviours is essential for interpreting the strengths and weaknesses of LLM-based investing strategies and for guiding their future development (Hui and Chan, 2018).
We label each calendar year based on the annual return of the S&P 500: $R_{y}=\frac{P_{T}-P_{0}}{P_{0}}$, where $P_{0}$ and $P_{T}$ are the adjusted closing prices on the first and last trading days of year $y$. A year is classified as bull if $R_{y}\geq+20\%$, bear if $R_{y}\leq-20\%$, and sideways otherwise. The $\pm 20\%$ threshold follows standard industry convention (Zweig, 2019).
@@ 8. Findings and Takeaways
First, we find that LLM-derived alpha is likely a methodological artefact of narrow, biased evaluations. The performance advantages reported in short-term, selective studies vanish under our bias-mitigated backtests, which reveal a consistent and statistically significant failure to generate alpha (Section 6.3). This suggests that current LLMs do not overcome the Efficient Market Hypothesis (Fama, 1970) in realistic conditions, and that prior gains stemmed from survivorship and look-ahead biases rather than genuine market inefficiency.
Second, model complexity does not equate to market competence. The scaling laws of natural language processing (Kaplan et al., 2020) do not translate effectively to financial markets, which impose intrinsic limits on extractable signals (Harvey and Liu, 2013). We show that larger models do not reliably outperform smaller ones, and both are consistently bettered by simpler models like ARIMA on risk-adjusted metrics (Table 4). Without encoded financial logic, architectural complexity appears to add noise rather than value.
Third, we diagnose “how” LLM agents fail, revealing a fundamental misalignment with market regimes. Our analysis (§7, Appendix F) shows that agents are pathologically miscalibrated: they are too conservative in bull markets and too aggressive in bear markets. This behavioural flaw contradicts the Adaptive Markets Hypothesis (Lo, 2004), shifting the issue from merely a lack of profitability to a more profound failure in the agents’ decision-making policies.

### v4

@@ 1. Introduction
Large language models (LLMs) are increasingly used in financial decision-making, especially for generating investment actions such as Buy, Hold, or Sell (ding2024largelanguagemodelagent; fatouros2025marketsenseai20enhancingstock). These so-called LLM timing-based investing strategies leverage LLMs’ ability to interpret historical and real-time data to autonomously trade. From sentiment-driven trading (zhang2024unveilingpotentialsentimentlarge) to sophisticated multi-agent systems (yu2023finmemperformanceenhancedllmtrading; finagentzhang2024multimodalfoundationagentfinancial), a growing body of work has explored the potential of LLMs as autonomous financial agents.
Backtesting is the standard method for assessing investment strategies, simulating them on historical data to evaluate profitability and robustness (chan2021quantitative). However, current LLM investing research suffers from fragmented, underdeveloped evaluation practices. Most studies assess performance over short periods, on few stock symbols, and often omit code release, limiting reproducibility. As summarised in Table 1, several recent methods evaluate over under a year, with fewer than ten stocks, and benchmark only against naïve baselines like Buy-and-Hold. Such short horizons and narrow stock universes lead to three well-documented sources of bias: survivorship bias (garcia1993survivorship), where delisted or failed stocks are omitted; look-ahead bias (chan2021quantitative), where future information inadvertently influences past decisions; and data-snooping bias (Bailey2015ThePO), where strategy performance is inflated through repeated testing on the same data. These biases can result in misleading performance assessments and undermine the validity of claimed improvements over traditional methods. This raises a central question: Can LLM-based investing strategies survive longer and broader robustness evaluations?
@@
While recent efforts such as wang2025quantbenchbenchmarkingaimethods and hu2025fintsbcomprehensivepracticalbenchmark have addressed benchmarking for deep learning (DL)-based trading and LLM-based time-series forecasting, comprehensive evaluation of LLM-based investing strategies remains unaddressed. Separately, FinBen (finben-2024) provides a thorough FinLLM benchmark covering multiple tasks, including decision-making. However, as a broad FinLLM benchmark, FinBen’s backtesting still relies on a limited, hand-picked symbol set, which contains the aforementioned biases and lacks a professional backtesting pipeline or systematic comparison with traditional strategies. To fill this gap, we introduce FINSABER, a comprehensive framework for benchmarking LLM timing-based investing strategies that supports longer backtesting periods, a broader and more diverse symbol universe, and explicit bias mitigation. Specifically, our main contributions are:
@@ 2. Related Works
Recent work using LLMs as investors directly employ LLMs to make investing decisions (ding2024largelanguagemodelagent). The most common approach leverages LLMs’ sentiment analysis capabilities, using either general-purpose LLMs (e.g., GPT, LLaMA (touvron2023llama), Qwen (bai2023qwentechnicalreport)) or fine-tuned financial variants like FinGPT (yang2023fingpt) to generate sentiment scores for trading decisions (lopezlira2023chatgptforecaststockprice; ruoxuportfolioperformance24; KIRTAC2024105227; zhang2024unveilingpotentialsentimentlarge). However, these approaches stop short of forming complete trading strategies, which require not only directional forecasts, but also realistic liquidity sizing for mitigating impact, development of execution rules for trade timing and risk management, and incorporation of trading costs.
More advanced approaches move beyond sentiment scores by summarising and reasoning over multi-source financial text. For example, fatouros2024largelanguagemodelsbeat introduce a memory module that stores summarised financial data, retrieved during trading to guide decisions. Similarly, LLMFactor (wang2024llmfactorextractingprofitablefactors) learns to extract profitable factors from historical news aligned with price movements and applies them to future market forecasts.
A growing body of work incorporates LLM-based agents (guo2024largelanguagemodelbased), where either one specialised agent or multiple collaborative agents are employed to perform financial analysis or predictions. Notable examples include FinMem (yu2023finmemperformanceenhancedllmtrading), FinAgent (finagentzhang2024multimodalfoundationagentfinancial), FinRobot (yang2024finrobotopensourceaiagent), TradExpert (ding2024tradexpertrevolutionizingtradingmixture), FinCon (yu2024fincon), TradingAgents (xiao2024tradingagents) and MarketSenseAI 2.0 (fatouros2025marketsenseai20enhancingstock). Some models also incorporate reinforcement learning (RL) for iterative self-improvement (RePEc:arx:papers:2310.05627; 10.1145/3589334.3645611).
@@ 4. Biases and Robustness Challenges in Backtesting LLM Investors
Robust evaluation of financial strategies demands carefully designed backtests. Unlike typical machine learning tasks with large, clean datasets, financial data is noisy, nonstationary, and limited in scope. As a result, backtests are especially prone to three major sources of bias: survivorship bias, look-ahead bias, and data-snooping bias, each of which can inflate perceived performance and lead to misleading conclusions (chan2021quantitative).
@@
This occurs when backtests include only currently active stocks while ignoring delisted or bankrupt assets. Such omissions systematically overstate returns and understate risk (joubert2024three). A common cause is using today’s S&P 500 constituents as the historical investment universe. This practice introduces what garcia1993survivorship call “preinclusion bias”, also a form of look-ahead bias where future index membership influences past decisions. The impact is well-documented: grinblatt1989mutual and elton1996survivor estimate annual return distortions between 0.1% and 0.9%, and brown1992survivorship show that even small distortions can misrepresent performance persistence.
@@
Look-ahead bias arises when a strategy uses information that would not have been known at the time of decision-making (chan2021quantitative). This includes selecting features, parameters, or symbols based on full-period outcomes, thereby introducing future knowledge into the backtest.
@@
Also known as multiple testing bias, this occurs when repeated experimentation on the same dataset leads to overfitting. In finance, where sample sizes are small and the signal-to-noise ratio is very low, this bias is particularly problematic. Bailey2015ThePO showed that evaluating strategies on overlapping data inflates false positive rates, and that standard hold-out validation techniques often fail to guard against this issue.
@@
Addressing these biases requires evaluating strategies across longer periods and broader asset universes. For daily trading, at least three years of data is generally recommended, while weekly and monthly strategies benefit from 10 to 20 years or more (Bailey2015ThePO). 10.1093/rfs/hhj020 tested pairs trading on 40 years of daily data, but do2010does extended this to 48 years and found profitability declined, highlighting the need for long-term evaluation. Likewise, recent deep learning models in finance rely on multi-year datasets to ensure robustness (fengtemporal; ijcai2022p0551; CI-STHPAN).
@@ 5. FINSABER
We incorporate a comprehensive collection of strategies across multiple paradigms to ensure robust benchmarking. The timing-based strategies include open-source LLM investors (FinMem (yu2023finmemperformanceenhancedllmtrading), FinAgent (finagentzhang2024multimodalfoundationagentfinancial)), traditional rule-based approaches (Buy and Hold, Moving Average Crossover, Bollinger Bands (bollinger2002bollinger), Trend Following (Wilcox2009trendfollowing)), ML/DL forecaster-based methods (ARIMA, XGBoost), and RL-based strategies (A2C, PPO, TD3, SAC implemented via FinRL (finrl-2022) framework). Selection-based strategies encompass random K selection, Momentum Factor Selection (based on past returns), Volatility Effect Selection (selecting low-volatility stocks), and the stocks selection agent from the FinCon (yu2024fincon) framework. This diverse strategy base enables comprehensive performance comparison across different methodological approaches while maintaining extensibility for custom implementations. More technical details of the strategies are available in Appendix B.
@@
High returns alone do not imply strategy quality. Risk-adjusted metrics such as SPR and STR are more informative, especially in finance where capital efficiency and downside risk are critical (chan2021quantitative). These metrics are standard in the literature (cont2001empirical; demiguel2009optimal) and widely used in recent LLM-based investing benchmarks (finagentzhang2024multimodalfoundationagentfinancial; yu2024fincon). Formal definitions and formulas are provided in Appendix C.
@@ 6.2. Fair Comparisons with the Composite Approach
To overcome the aforementioned biases, we introduce the Composite evaluation setup within FINSABER. This setup integrates systematic selection-based strategies to expand and diversify the stock universe, explicitly addressing survivorship and data-snooping biases. Specifically, we use four unbiased stock selection approaches from the strategies base (details in Appendix B): Random Five, Momentum Factor (Muller31032010), Volatility Effect (volatility-effect), and the FinCon Selection Agent in the FinCon (yu2024fincon) framework.
@@ 6.3. Statistical Validation and Behavioural Diagnostics of LLM Agents
To validate our findings from the composite backtests and diagnose the underlying drivers of LLM agent performance, we conduct a unified statistical and behavioural analysis. First, we conduct paired t-tests comparing Buy and Hold, FinMem, and FinAgent across both Selected 4 (Table 3) and Composite (Table 4) setups. Second, we dissect the agents’ behavioural characteristics by examining their drawdown profiles, alpha ($\alpha$) and beta ($\beta$) decomposition, and trading turnover across the different selection environments. These metrics are obtained by regressing the strategy’s excess returns against the market’s excess returns based on the Capital Asset Pricing Model (CAPM) (capm-1964). The model is defined as: $R_{s}-R_{f}=\alpha+\beta(R_{m}-R_{f})+\epsilon$, where $R_{s}$ is the return of the strategy, $R_{m}$ is the market return, $R_{f}$ is the risk-free rate, and $\epsilon$ is the idiosyncratic residual. In this model, $\beta$ measures the strategy’s systematic risk or volatility relative to the market, while $\alpha$ represents the portion of the return not explained by market exposure, often considered a measure of strategy-specific skill.
@@
Table 5 reports t-tests and p-values for the previous results, testing the null hypothesis of equal performance distributions. Under the selective period, statistical significance is inconsistent and limited mostly to individual stocks. However, after mitigating biases through the composite setup, the p-values drop substantially, indicating the market baseline (Buy and Hold) significantly outperforms both LLM strategies across all robust setups. Notably, while FinAgent tends to outperform FinMem when biases are controlled, both still significantly underperform simple market baselines. Furthermore, the behavioural analysis in Table 6 reveals that this underperformance is rooted in a lack of genuine skill; neither LLM agent generates statistically significant alpha, with all measured p-values exceeding 0.34. This finding robustly supports our main thesis that the claimed superiority of these models does not hold under rigorous evaluation, aligning with the Efficient Market Hypothesis (adaptive-markets-hypo).
@@ 7. Market Regime Analysis
Another key question in evaluating LLM-based investing strategies is whether they adapt appropriately across varying market conditions. Financial markets exhibit time-varying predictability and uncertainty across different economic, financial, and political regimes (kim2011stock). Some strategies may exploit these variations, while others may struggle to adapt. Distinct market environments—bull, bear, and sideways—present unique challenges and opportunities: bull markets reward aggressive positioning and high exposure, bear markets require effective risk management, and sideways markets test a strategy’s ability to navigate uncertainty in the absence of clear trends. By decomposing performance across these regimes, it is possible to determine whether strategies are overly conservative and miss opportunities during bullish periods, or excessively aggressive and incur significant losses during downturns. Understanding these regime-specific behaviours is essential for interpreting the strengths and weaknesses of LLM-based investing strategies and for guiding their future development (optimaltradingstrategyduringbullbear).
We label each calendar year based on the annual return of the S&P 500: $R_{y}=\frac{P_{T}-P_{0}}{P_{0}}$, where $P_{0}$ and $P_{T}$ are the adjusted closing prices on the first and last trading days of year $y$. A year is classified as bull if $R_{y}\geq+20\%$, bear if $R_{y}\leq-20\%$, and sideways otherwise. The $\pm 20\%$ threshold follows standard industry convention (wallstreetbullbear).
@@ 8. Findings and Takeaways
First, we find that LLM-derived alpha is likely a methodological artefact of narrow, biased evaluations. The performance advantages reported in short-term, selective studies vanish under our bias-mitigated backtests, which reveal a consistent and statistically significant failure to generate alpha (Section 6.3). This suggests that current LLMs do not overcome the Efficient Market Hypothesis (fama1970efficient) in realistic conditions, and that prior gains stemmed from survivorship and look-ahead biases rather than genuine market inefficiency.
Second, model complexity does not equate to market competence. The scaling laws of natural language processing (kaplan2020scalinglawsneurallanguage) do not translate effectively to financial markets, which impose intrinsic limits on extractable signals (backtesting-ssrn). We show that larger models do not reliably outperform smaller ones, and both are consistently bettered by simpler models like ARIMA on risk-adjusted metrics (Table 4). Without encoded financial logic, architectural complexity appears to add noise rather than value.
Third, we diagnose “how” LLM agents fail, revealing a fundamental misalignment with market regimes. Our analysis (§7, Appendix F) shows that agents are pathologically miscalibrated: they are too conservative in bull markets and too aggressive in bear markets. This behavioural flaw contradicts the Adaptive Markets Hypothesis (adaptive-markets-hypo), shifting the issue from merely a lack of profitability to a more profound failure in the agents’ decision-making policies.